if sin(50degree – 3/2a) = cos(3 alpha – 50degree ) , then findthe value of alpha and hence evaluate tan alpha.secalpha.sinalpha – cotalpha.sinalpha.cosalpha 

Grade:12

1 Answers

Shubham Kamleshbhai Patel
8 years ago
This is a easy one.
U need to use here cos a = sin ( 90 - a )
 
sin ( 50° – 3/2 a ) = cos ( 3a – 50° )
=> sin ( 50° – 3/2a ) = sin ( 120° – 3a )
 
U will the value of a = 60°
 
Thus, applying a in tan a x sec acot a x sin a x cos a
=> root 3 x 2 – 1/root 3 x root3 / 2 x ½
=> 4 root 3 – 1 / 4
 
Do check the answer and reply back......